Sin A +SinB =root 3(cosA-cosB) sin3A -sin 3B is equal to

Grade:9

1 Answers

Arun
5 years ago
sinA + sinB = root(3) cosA - root(3)cosB

Divide by 2 both side

sinA / 2  + sinB / 2  = root(3).cosA/2 - root(3).cosB/2

1/2 . sinA  - root(3)/2 .cosA  = - root(3)/2 . cos(B)  - 1/2 . sin(B)

sin(A - 60)  = - sin(B + 60) 

so finaly  A -60= - 60 - B

so A = -B

Therefore, Sin3B + sin3A = 0